OpenStudy (winsteria):

MEDAL AND FAN Maura had a pizza delivered to her home. The total amount she paid was $17.53. Included in the total amount were the menu price of the pizza, 6% sales tax on the price of the pizza, and a $3.75 delivery fee. What was the menu price of the pizza?

OpenStudy (winsteria):

@sammixboo

OpenStudy (winsteria):

No

OpenStudy (winsteria):

okay

OpenStudy (winsteria):

@janeSTAR

OpenStudy (winsteria):

help me with it?

OpenStudy (anonymous):

Okay let's take this backwards for a moment. Say you buy something at Wal-Mart for $10 (before tax). When you get to the register, you will have to pay tax of 6%. To find your total price, it will be the $10*(100 + 6)% = $10*1.06 = $10.60

OpenStudy (winsteria):

okay

OpenStudy (anonymous):

Since the delivery charge of the pizza isn't taxed, it's simply added to the subtotal price. So in the example I made up, if you were to say it costs $5 in gas to get home, you'd simply add that to the $10.60 to get $15.60. To solve your problem, we're simply doing this in reverse.

OpenStudy (winsteria):

okay...

OpenStudy (anonymous):

Did that help any?

OpenStudy (winsteria):

I think It's 13 but I'm not sure @janeSTAR

OpenStudy (anonymous):

Okay I think its correct .How did I solve it : 17.53 - 3.75 = 13.78 = T P = T/1.06 = 13 So the cost on the menu, before tax and delivery, is $13

OpenStudy (winsteria):

Thank you

OpenStudy (anonymous):

NP